Simplify the following expression as much as possible:
3 square root of 48 x square root of 12

Answer by fractalier(6550)

3*sqrt(48) * sqrt(12) =
3*sqrt(16)*sqrt(3) * sqrt(4)*sqrt(3) =
12*sqrt(3) * 2*sqrt(3) =
24 * 3 =
72